# Adaptive and Efficient Learning with Blockwise Missing and Semi-Supervised Data

### v1

@@ 1.2 Problem Setup
The model parameter $\boldsymbol{\gamma}=(\gamma_{1},\ldots,\gamma_{p})^{\top}$ can be identified and estimated using the standard regression on the $\mathcal{L}\mathcal{C}$ sample but $\mathcal{L}\mathcal{M}$ and $\mathcal{U}\mathcal{C}$ cannot be directly incorporated into this procedure, which results in loss of effective samples. In this paper, our aim is to leverage both the $\mathcal{L}\mathcal{M}$ and $\mathcal{U}\mathcal{C}$ samples to enhance the estimation efficiency of $\boldsymbol{\gamma}$ without incurring bias.
@@ 1.3 Related Literature
As will be detailed in our asymptotic and numerical studies, the above-introduced reduced model or imputation strategies does not characterize the true influence of the $\mathcal{L}\mathcal{M}$ data and, thus, not achieving semiparametric efficiency in the BM regression setting, which dates back to the seminal work by Robins et al., (1994) and has been explored (Chen et al.,, 2008; Li and Luedtke,, 2023; Li et al., 2023a, ; Qiu et al.,, 2023, e.g.) in various settings. In particular, Li and Luedtke, (2023) proposed derived the canonical gradient for a general data fusion problem and justified its efficient. Nevertheless, such semiparametric efficiency property is established under an ideal scenario with a correct conditional model for $\boldsymbol{X}_{\mathcal{P}_{m}}$, which is stringent in practice. Their method could be less robust and effective in a more realistic case when the imputation for $\boldsymbol{X}_{\mathcal{P}_{m}}$ is not consistent with the truth. This issue has been recognized and addressed in recent literature mainly focusing on the SS inference problem (Gronsbell et al.,, 2022; Schmutz et al.,, 2022; Angelopoulos et al.,, 2023; Miao et al.,, 2023; Gan and Liang,, 2023; Wu and Liu,, 2023).
@@ 1.4 Our contribution
    We develop a novel data-driven control function approach to realize more adaptive and efficient estimation under potential misspecification or poor quality of the imputation models, especially when there exists nonlinear relationship not captured by them. Compared to recent work addressing problematic nuisance models in a more simple SS scenario (Angelopoulos et al.,, 2023; Miao et al.,, 2023, e.g.), our strategy can be generalized to their settings and realize more flexible and efficient calibration.
@@
  4. 4.
    We establish a novel strategy to bring in large unlabeled samples under the SS setting and leverage them to enhance the efficiency of the BM estimation. This enables a more efficient use of the $\mathcal{U}\mathcal{C}$ sample leading to an estimator of strictly smaller variance than that obtained only using $\mathcal{L}\mathcal{C}$ and $\mathcal{L}\mathcal{M}$, which has not been readily achieved in previous work on a similar setting (Xue et al.,, 2021; Song et al.,, 2024). Our development is also shown to be more efficient than existing approaches for the standard SS setting without BM (Gronsbell et al.,, 2022, e.g.), in both theoretical and numerical studies.
@@ 2.1 Outline and Preliminary Framework
The data imputation model $\tilde{{\mu}}_{\boldsymbol{X}_{\mathcal{P}_{m}}}$ can be either a simple parametric model with finite-dimensional $\tilde{\boldsymbol{\eta}}$ or more sophisticated ones with diverging parameter space. For example, for continuous $\boldsymbol{X}_{\mathcal{P}_{m}}$, one can either (i) assume that $\boldsymbol{X}_{\mathcal{P}_{m}}$ follows a (multivariate) Gaussian linear model against $\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y$ and learn it through the a standard MLE; or (ii) assume $\boldsymbol{X}_{\mathcal{P}_{m}}=m(\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y)+\boldsymbol{\varepsilon}$ with $\boldsymbol{\varepsilon}\sim{\rm N}(0,\boldsymbol{\Sigma})$ independent from $\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y$, then use some ML algorithm (e.g., random forest or neural network) to learn the mean model $m(\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y)$ and the standard method of moment on the fitted residuals to estimate $\boldsymbol{\Sigma}$; or (iii) implement generative ML like conditional generative adversarial network (GAN) (Mirza and Osindero,, 2014) to learn ${\mu}$ without any model assumptions. Regarding these choices, our asymptotic analysis generally accommodate all the three types, and we use both (i) and (ii) in our simulation and real-world studies.
@@
When using complex ML algorithms like random forests to learn $\tilde{{\mu}}_{\boldsymbol{X}_{\mathcal{P}_{m}}}$, we adopt cross-fitting to avoid the over-fitting bias as frequently done in related contexts (Chernozhukov et al.,, 2016; Liu et al.,, 2023, e.g.). In specific, we randomly split the $\mathcal{L}\mathcal{C}$ sample into $K$ folds with an equal size, denoted as $\{\mathcal{L}\mathcal{C}^{(1)},\mathcal{L}\mathcal{C}^{(2)},...,\mathcal{L}\mathcal{C}^{(K)}\}$. For each $k\in[K]$, we implement the ML method on $\mathcal{L}\mathcal{C}^{(-k)}=\cup_{j\neq k}\mathcal{L}\mathcal{C}^{(j)}$ to obtain an estimator $\tilde{\boldsymbol{\eta}}^{(-k)}$. Then for each $i$ from $\mathcal{L}\mathcal{C}$, we use $\tilde{\varphi}_{1}(\boldsymbol{X}_{i,\mathcal{P}_{m^{c}}},Y_{i};\tilde{\boldsymbol{\eta}}^{(-k)})$ as its score imputation since $\tilde{\boldsymbol{\eta}}^{(-k)}$ is independent with $\boldsymbol{D}_{i}$. For each $i$ from $\mathcal{L}\mathcal{M}$, we use the average imputation $K^{-1}\sum_{k=1}^{K}\tilde{\varphi}_{1}(\boldsymbol{X}_{i,\mathcal{P}_{m^{c}}},Y_{i};\tilde{\boldsymbol{\eta}}^{(-k)})$. One can find more details about this cross-fitting procedure in Appendix. As shown in Section 3, this strategy is important to eliminate the over-fitting bias of arbitrary ML methods. For the ease of presentation, we moderately abuse the notation in the remaining paper by dropping the superscript of $k$ related to cross-fitting.
@@
With $\tilde{{\mu}}_{\boldsymbol{X}_{\mathcal{P}_{m}}}$, one can use either numerical methods like Gaussian quadrature or Monte-Carlo (MC) sampling to compute $\tilde{\varphi}_{1,i}$. When ${\boldsymbol{X}_{\mathcal{P}_{m}}}$ is multiple or high dimensional, the computational burden and accuracy through either way seems to be a challenging issue. However, inspecting the form of the score function $\boldsymbol{c}^{\top}\tilde{\boldsymbol{S}}(\boldsymbol{X},Y)$, we notice that it depends on the entire $\boldsymbol{X}_{\mathcal{P}_{m}}$ only through at most two-dimensional linear representations $\tilde{\boldsymbol{\gamma}}_{\mathcal{P}_{m}}^{\top}\boldsymbol{X}_{\mathcal{P}_{m}}$ and $(\boldsymbol{c}^{\top}\bar{\boldsymbol{H}}^{-1})_{\mathcal{P}_{m}}\boldsymbol{X}_{\mathcal{P}_{m}}$. Thus, both Gaussian quadrature and MC can achieve good computational performance thanks to this low dimensionality (Caflisch,, 1998; Kovvali,, 2022). For MC, we could also use the same set of MC samples on all subjects in certain scenarios (e.g., Gaussian noise under homoscedasticity) to save computational costs. In our case, we choose the MC method to compute $\tilde{\varphi}_{1}$ as it facilitates our data-adaption procedure to be introduced next.
@@ 2.2 Data-adaptive Calibration
Following the definition of $w$, we can construct $\mu^{\prime}$ as the product of $\mu$ and an IW function $w$ that is only parameterized by $\boldsymbol{\delta}$. It is easy to see that when $w\equiv 1$, $\mu^{\prime}\equiv\mu$. Hence, $\mu^{\prime}$ encompasses a broader class of conditional models of $\boldsymbol{X}_{P_{m}}$. We then replace $\varphi^{*}$ in (5) by $\tilde{\varphi}_{1}^{\prime}$, and calibrate it by optimizing $\boldsymbol{\delta}$ as follows.
@@ 2.3 Semi-supervised Estimation
$\hat{\beta}_{2}=\hat{\beta}_{1}-\widehat{{\rm E}}_{\mathcal{L}\mathcal{C}}[\tilde{\varphi}_{2}^{\prime}(\boldsymbol{X};\boldsymbol{\zeta},\tilde{\boldsymbol{\theta}})]+\widehat{{\rm E}}_{\mathcal{U}\mathcal{C}}[\tilde{\varphi}_{2}^{\prime}(\boldsymbol{X};\boldsymbol{\zeta},\tilde{\boldsymbol{\theta}})].$ (9)
@@ 2.4 DEFUSE in the Multiple BM Scenario
2.5 Extension on Auxiliary Covariates
We shall end the method section with brief description and discussion on the extension of DEFUSE to accommodate the situation with auxiliary features frequently encountered in many real-world applications. Consider the scenario with $\boldsymbol{X}=\{\boldsymbol{A},\boldsymbol{W}\}$ consisting of some primary risk factors $\boldsymbol{A}$ and auxiliary or surrogate features $\boldsymbol{W}$. Instead of the outcome model against the whole $\boldsymbol{X}$, now the focus is on $Y\sim\boldsymbol{A}$ with the population equation ${\rm E}\left[\boldsymbol{A}\left\{Y-g\left(\boldsymbol{\gamma}^{\top}\boldsymbol{A}\right)\right\}\right]=\mathbf{0}$. Meanwhile, the missing structure remains the same as described in Section 1.2, i.e., one or multiple subsets of covariates $\boldsymbol{X}_{\mathcal{P}_{m}}$ partially missing on the labeled sample and $Y$ being unlabeled on $\mathcal{U}\mathcal{C}$ with the complete $\boldsymbol{X}$. Here, covariates $\boldsymbol{A}$ can be either included by, intersected with, or disjoint with $\boldsymbol{X}_{\mathcal{P}_{m}}$.
This is a common setup in real-world studies. For example, on EMR linked biobank data, the interest lies in the relationship between certain genetic variants $\boldsymbol{A}$ and disease status $Y$. $\boldsymbol{W}$ is taken as EMR proxies (e.g., relevant codified features) for $Y$ and acts as auxiliary features not included in the genetic risk model. In clinical or epidemiological studies, $\boldsymbol{A}$ is the treatment or key risk factors, $Y$ is a long-term outcome, and $\boldsymbol{W}$ can be some early endpoints or surrogate outcomes (e.g., tumor response rates) occurring and observed post the baseline with only $\boldsymbol{A}$. In this case, one is interested in $Y\sim\boldsymbol{A}$ to make clinical decision at the baseline without $\boldsymbol{W}$. In both examples, the auxiliary $\boldsymbol{W}$ could still be informative to $Y$ and, thus, included to boost the statistical efficiency.
Our framework can be naturally and easily extended to accommodate the setting introduced above. In specific, the only change to make is to replace the Step 1 in Algorithm 1 with the one solving for $\tilde{\boldsymbol{{\gamma}}}$ from $\widehat{{\rm E}}_{\mathcal{L}\mathcal{C}}[\boldsymbol{A}\{Y-g(\boldsymbol{\gamma}^{\top}\boldsymbol{A})\}]=\mathbf{0}$ and deriving the corresponding score function $\tilde{\boldsymbol{S}}_{i}=\tilde{\boldsymbol{H}}^{-1}\boldsymbol{A}_{i}\{Y_{i}-g(\tilde{\boldsymbol{{\gamma}}}^{\top}\boldsymbol{A}_{i})\}$ where $\tilde{\boldsymbol{H}}=\widehat{{\rm E}}_{\mathcal{L}\mathcal{C}}[g(\tilde{\boldsymbol{{\gamma}}}^{\top}\boldsymbol{A})\boldsymbol{A}\boldsymbol{A}^{\top}]$. All the conditional modeling, imputation, and calibration procedures in the following Steps 2 and 3 in Algorithm 1 can be conducted in exactly the same forms as presented in previous sections. In this case, $\boldsymbol{W}$ could often be of high-dimensionality as considered in Zhou et al., (2024). Our theoretical framework to be established in the next section can also be applied to this setting.
@@ 3.1 Notations and Assumptions
$\Delta_{\mu}(\tilde{\boldsymbol{\eta}},\bar{\boldsymbol{\eta}})=\left|1-\frac{{\mu}(\boldsymbol{X}_{\mathcal{P}_{m}}|\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y;\tilde{\boldsymbol{\eta}})}{{\mu}(\boldsymbol{X}_{\mathcal{P}_{m}}|\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y;\bar{\boldsymbol{\eta}})}\right|;\quad\Delta_{\nu}(\tilde{\boldsymbol{\theta}},\bar{\boldsymbol{\theta}})=\left|1-\frac{\nu(Y|\boldsymbol{X};\tilde{\boldsymbol{\theta}})}{\nu(Y|\boldsymbol{X};\bar{\boldsymbol{\theta}})}\right|;$
@@
${\rm Cov}\left[\bar{\boldsymbol{\varphi}}_{1}^{\bullet}(\boldsymbol{X},Y)\right]+{\rm Diag}\left\{\frac{1}{\bar{\rho}_{r}}{\rm Var}\left[\bar{\varphi}_{1,r}^{\prime}(\boldsymbol{X}_{\mathcal{P}_{m^{c}}^{1}},Y)\right]\right\}_{r=1,\ldots,R}$

### v2

@@ 1.2 Problem Setup
The model parameter $\boldsymbol{\gamma}=(\gamma_{1},\ldots,\gamma_{p})^{\top}$ can be identified and estimated using the standard regression on the $\mathcal{L}\mathcal{C}$ sample but $\mathcal{L}\mathcal{M}$ and $\mathcal{U}\mathcal{C}$ cannot be directly incorporated into this procedure, which results in loss of effective samples. In this paper, our aim is to leverage both the $\mathcal{L}\mathcal{M}$ and $\mathcal{U}\mathcal{C}$ samples to enhance the estimation efficiency of $\boldsymbol{\gamma}$ without incurring bias. As will be introduced in Section 5.1, the data in our setup can also include auxiliary features that are predictive of $Y$ and $\boldsymbol{X}$, observed on all subjects, and possibly high-dimensional. These variables are not included in the outcome model $Y\sim\boldsymbol{X}$ of our primary interest due to scientific reasons but can be simply incorporated into our framework to enhance estimation efficiency. For the ease of presentation, we drop the notation of such auxiliary features in the main body of our method and theory.
@@ 1.3 Related Literature
As will be detailed in our asymptotic and numerical studies, the above-introduced reduced model or imputation strategies does not characterize the true influence of the $\mathcal{L}\mathcal{M}$ data and, thus, not achieving semiparametric efficiency in the BM regression setting, which dates back to the seminal work by Robins et al., (1994) and has been explored (Chen et al.,, 2008; Li and Luedtke,, 2023; Li et al., 2023a, ; Qiu et al.,, 2023, e.g.) in various settings. In particular, Li and Luedtke, (2023) proposed derived the canonical gradient for a general data fusion problem and justified its efficient. Nevertheless, such semiparametric efficiency property is established under an ideal scenario with a correct conditional model for $\boldsymbol{X}_{\mathcal{P}_{m}}$, which is stringent in practice. Their method could be less robust and effective in a more realistic case when the imputation for $\boldsymbol{X}_{\mathcal{P}_{m}}$ is not consistent with the truth. This issue has been recognized and studied in recent literature mainly focusing on SS inference (Gronsbell et al.,, 2022; Schmutz et al.,, 2022; Song et al.,, 2023; Angelopoulos et al.,, 2023; Miao et al.,, 2023; Gan and Liang,, 2023; Wu and Liu,, 2023). Notably, in our real-world study including high-dimensional auxiliary features to impute the missing variables (see Section 5.1), this problem is even more pronounced due to the difficulty of high-dimensional modeling. In this case, it is challenging to realize adaptive and efficient data fusion with above-mentioned strategies, which can be seen from the results in Section 5.
@@ 1.4 Our contribution
    We develop a novel data-driven control function approach to realize adaptive and efficient estimation under potential misspecification or poor quality of the imputation models, especially when there exists nonlinear relationship not captured by them. Compared to recent work in addressing problematic nuisance models (Angelopoulos et al.,, 2023; Miao et al.,, 2023, e.g.), our strategy is shown to realize more flexible and efficient calibration in our theory, simulation, and real-world studies. This improvement is even clearer when high-dimensional auxiliary features are included in the nuisance model like our EMR application.
@@
    We establish a novel framework to bring in large unlabeled samples under the SS setting and leverage them to enhance the efficiency of the BM estimation. This enables a more efficient use of the $\mathcal{U}\mathcal{C}$ sample leading to an estimator of strictly smaller variance than that obtained only using $\mathcal{L}\mathcal{C}$ and $\mathcal{L}\mathcal{M}$, which has not been readily achieved in previous work on a similar setting (Xue et al.,, 2021; Song et al.,, 2024). Our development is also shown to be more efficient than existing approaches for the standard SS setting without BM (Gronsbell et al.,, 2022, e.g.), in both theoretical and numerical studies.
  4. 4.
@@ 2.1 Outline and Preliminary Framework
The data imputation model $\tilde{{\mu}}_{\boldsymbol{X}_{\mathcal{P}_{m}}}$ can be either a simple parametric model with finite-dimensional $\tilde{\boldsymbol{\eta}}$ or more sophisticated ones with diverging parameter space. For example, for continuous $\boldsymbol{X}_{\mathcal{P}_{m}}$, one can either (i) assume that $\boldsymbol{X}_{\mathcal{P}_{m}}$ follows a (multivariate) Gaussian linear model against $\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y$ and learn it through the a standard MLE; or (ii) assume $\boldsymbol{X}_{\mathcal{P}_{m}}=m(\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y)+\boldsymbol{\varepsilon}$ with $\boldsymbol{\varepsilon}\sim{\rm N}(0,\boldsymbol{\Sigma})$ independent from $\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y$, then use some ML algorithm (e.g., random forest or neural network) to learn the mean model $m(\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y)$ and the standard method of moment on the fitted residuals to estimate $\boldsymbol{\Sigma}$; or (iii) implement generative ML like conditional generative adversarial network (GAN) (Mirza and Osindero,, 2014) to learn ${\mu}$ without any model assumptions.
Moreover, as will be discussed and implemented in Section 5, $\tilde{{\mu}}_{\boldsymbol{X}_{\mathcal{P}_{m}}}$ can be constructed with auxiliary covariates not belonging to $\boldsymbol{X}$ included in the targeted outcome model. This may invoke learning with high-dimensional features and excessive errors, which highlights the importance of carrying out flexible data-adaptive calibration to be introduced in Section 2.2. Regarding these choices, our asymptotic analysis generally accommodate all the three types with or without auxiliary covariates.
@@
When using complex ML algorithms like random forests to learn $\tilde{{\mu}}_{\boldsymbol{X}_{\mathcal{P}_{m}}}$, we adopt cross-fitting to avoid the over-fitting bias as frequently done in related contexts (Chernozhukov et al.,, 2016; Liu et al.,, 2023, e.g.). In specific, we randomly split the $\mathcal{L}\mathcal{C}$ sample into $K$ folds with an equal size, denoted as $\{\mathcal{L}\mathcal{C}^{(1)},\mathcal{L}\mathcal{C}^{(2)},...,\mathcal{L}\mathcal{C}^{(K)}\}$. For each $k\in[K]$, we implement the ML method on $\mathcal{L}\mathcal{C}^{(-k)}=\cup_{j\neq k}\mathcal{L}\mathcal{C}^{(j)}$ to obtain an estimator $\tilde{\boldsymbol{\eta}}^{(-k)}$. Then for each $i$ from $\mathcal{L}\mathcal{C}$, we use $\tilde{\varphi}_{1}(\boldsymbol{X}_{i,\mathcal{P}_{m^{c}}},Y_{i};\tilde{\boldsymbol{\eta}}^{(-k)})$ as its score imputation since $\tilde{\boldsymbol{\eta}}^{(-k)}$ is independent with $\boldsymbol{D}_{i}$. For each $i$ from $\mathcal{L}\mathcal{M}$, we use the average imputation $K^{-1}\sum_{k=1}^{K}\tilde{\varphi}_{1}(\boldsymbol{X}_{i,\mathcal{P}_{m^{c}}},Y_{i};\tilde{\boldsymbol{\eta}}^{(-k)})$. As shown in Section 3, this strategy is important to eliminate the over-fitting bias of arbitrary ML methods. For the ease of presentation, we slightly abuse the notation in the remaining paper by dropping the superscript of $k$ related to cross-fitting.
@@
With $\tilde{{\mu}}_{\boldsymbol{X}_{\mathcal{P}_{m}}}$, one can use either numerical methods like Gaussian quadrature or Monte-Carlo (MC) sampling to compute $\tilde{\varphi}_{1,i}$. When ${\boldsymbol{X}_{\mathcal{P}_{m}}}$ includes many covariates, the computational burden and accuracy through either way seems to be a challenging issue. However, inspecting the form of the score function $\boldsymbol{c}^{\top}\tilde{\boldsymbol{S}}(\boldsymbol{X},Y)$, we notice that it depends on the entire $\boldsymbol{X}_{\mathcal{P}_{m}}$ only through at most two-dimensional linear representations $\tilde{\boldsymbol{\gamma}}_{\mathcal{P}_{m}}^{\top}\boldsymbol{X}_{\mathcal{P}_{m}}$ and $(\boldsymbol{c}^{\top}\bar{\boldsymbol{H}}^{-1})_{\mathcal{P}_{m}}\boldsymbol{X}_{\mathcal{P}_{m}}$. Thus, both Gaussian quadrature and MC can achieve good computational performance thanks to this low dimensionality (Caflisch,, 1998; Kovvali,, 2022). For MC, we could also use the same set of MC samples on all subjects in certain scenarios (e.g., Gaussian noise under homoscedasticity) to save computational costs. In our case, we choose the MC method to compute $\tilde{\varphi}_{1}$ as it facilitates our data-adaption procedure to be introduced next.
@@ 2.2 Data-adaptive Calibration
Following the definition of $w$, we can construct $\mu^{\prime}$ as the product of $\mu$ and an IW function $w$ that is only parameterized by $\boldsymbol{\delta}$. It is easy to see that when $w\equiv 1$, $\mu^{\prime}\equiv\mu$. Hence, $\mu^{\prime}$ encompasses a broader class of conditional models of $\boldsymbol{X}_{\mathcal{P}_{m}}$. We then replace $\varphi^{*}$ in (5) by $\tilde{\varphi}_{1}^{\prime}$, and calibrate it by optimizing $\boldsymbol{\delta}$ as follows.
@@ 2.3 Semi-supervised Estimation
$\hat{\beta}_{2}=\hat{\beta}_{1}-\widehat{{\rm E}}_{\mathcal{L}\mathcal{C}}[\tilde{\varphi}_{2}^{\prime}(\boldsymbol{X};\tilde{\boldsymbol{\zeta}},\tilde{\boldsymbol{\theta}})]+\widehat{{\rm E}}_{\mathcal{U}\mathcal{C}}[\tilde{\varphi}_{2}^{\prime}(\boldsymbol{X};\tilde{\boldsymbol{\zeta}},\tilde{\boldsymbol{\theta}})].$ (9)
@@ 3.1 Notations and Assumptions
$\Delta_{\mu}(\tilde{\boldsymbol{\eta}},\bar{\boldsymbol{\eta}})=\left|{\mu}(\boldsymbol{X}_{\mathcal{P}_{m}}|\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y;\tilde{\boldsymbol{\eta}})-{\mu}(\boldsymbol{X}_{\mathcal{P}_{m}}|\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y;\bar{\boldsymbol{\eta}})\right|;$
$\Delta_{\nu}(\tilde{\boldsymbol{\theta}},\bar{\boldsymbol{\theta}})=\left|\nu(Y|\boldsymbol{X};\tilde{\boldsymbol{\theta}})-\nu(Y|\boldsymbol{X};\bar{\boldsymbol{\theta}})\right|;$
@@
${\rm Var}\left[\bar{\boldsymbol{\varphi}}_{1}^{\bullet}(\boldsymbol{X},Y)\right]+{\rm Diag}\left\{\frac{1}{\bar{\rho}_{r}}{\rm Var}\left[\bar{\varphi}_{1,r}^{\prime}(\boldsymbol{X}_{\mathcal{P}_{m^{c}}^{r}},Y)\right]\right\}_{r=1,\ldots,R}$